# Effective Polaron Dynamics of an Impurity Particle Interacting with a Fermi Gas

### v1

@@ Abstract
We study the quantum dynamics of a homogeneous ideal Fermi gas coupled to an impurity particle on a three-dimensional box with periodic boundary condition. For large Fermi momentum $k_{\textnormal{F}}$, we prove that the effective dynamics is generated by a Fröhlich-type polaron Hamiltonian, which linearly couples the impurity particle to an almost-bosonic excitation field. Moreover, we prove that the effective dynamics can be approximated by an explicit coupled coherent state. Our method is applicable to two relevant settings: first, an interaction coupling $\lambda=1$ and masses of order 1 for time scales of order $k_{\textnormal{F}}^{-1}$; second to the case of $\lambda=k_{\textnormal{F}}^{-1}$ and a heavy Fermi gas with masses of order $k_{\textnormal{F}}$ for time scales of order 1.
@@ 1 Introduction
The study of impurities in quantum gases has garnered considerable attention due to its relevance in various physical contexts, ranging from solid-state physics to cold atom experiments. In this context quasi-particles such as polarons stand as intriguing entities emerging from the interaction of a single impurity particle with a surrounding medium. The concept of a polaron, originally introduced by Lev Landau to study the motion of an electron in a dielectric crystal [1965], most famously emerges from the celebrated Fröhlich Hamiltonian in second quantization formalism describing electron-phonon interactions [Froehlich1954]. Subsequently, the polaron concept was extended to all kind of surrounding media including Bose and Fermi gases. The formation conditions and properties of polarons are believed to play a central role to understand the transport properties and the effective mass of impurities within the host material.
In this article, we study with mathematical rigor the dynamics of an impurity particle immersed in a dense gas of fermions as surrounding medium. Interactions between fermions are neglected and we assume that the initial state $\psi$ of the system is a product state between the impurity state and a filled Fermi ball. This mathematical framework finds resonance with recent experimental and theoretical advancements in the study of ultracold atoms [SWSZ09, KSN+12, CJL+15]. We show that the effective dynamics of the system is governed by a Fröhlich-type Hamiltonian, which linearly couples the impurity particle to an almost-bosonic excitation field. More specifically, the excitations relative to the filled Fermi ball are up to a constant described by the Hamiltonian
@@
Subsequently, we show that the effective time evolved state can be up to a phase factor approximated by a time-dependent coupled coherent state $W(\eta_{t})\phi\otimes\Omega$ where $W$ is the Weyl operator which is simply parameterized by a function $\eta_{t}$. An explicit expression for $\eta_{t}$ is derived which allows for determining the number of collective excitations over time. We believe that such quantities are in particularly helpful to gain deeper insights into the formation process of quasi-particles as studied in experiments such as [CJL+16]. Eventually, we show that the linear coupling term $\Phi(h_{y})$ cannot be omitted in the effective description but adds a leading order effect to the effective dynamics in our setting.
Our results hold for a variety of time scales and couplings, describing different interaction strengths and mass ratio, which will be specified in the subsequent section. We note that the same microscopic model has been studied in [JMPP17, JMP18, MP21] but with very specific choices of couplings different from ours, leading to an effective decoupling of impurity and gas.
@@ 1.1 The microscopic model
We consider an impurity particle interacting with $N$ spinless fermions on a 3-dimensional box with periodic boundaries described by $\Lambda\coloneqq\mathbb{T}^{3}\coloneqq\mathbb{R}^{3}/(2\pi\mathbb{Z}^{3})$. The system is described by a state in the Hilbert space $L^{2}(\Lambda,\textnormal{d}y)\otimes\mathcal{H}_{N}^{-}$ with $\mathcal{H}_{N}^{-}=L^{2}(\Lambda)^{\wedge N}$ where $y$ is the coordinate of the impurity particle and $\{x_{i}\}_{i=1,\ldots,N}$ are the coordinates of the fermions. The Hamiltonian of the system is given by
$H_{N}=-\beta\Delta_{y}+\kappa\sum_{i=1}^{N}(-\Delta_{x_{i}})+\lambda\sum_{i=1}^{N}V(x_{i}-y)$ (1.2)
and parameterized by the coupling constants $\beta,\kappa,\lambda>0$. Note that the different parts of the Hamiltonian on different tensor components of our Hilbert space writing, i.e. we used the short-hand notation writing, for example, $-\Delta_{y}\equiv-\Delta_{y}\otimes 1$ for the Laplacian acting on the impurity particle. The interaction $V$ is assumed to have a Fourier transform $\hat{V}$ with compact support satisfying $\hat{V}(k)=\hat{V}(-k)$ for all $k\in\mathbb{Z}^{3}$. It is well-known that under this assumption the Hamiltonian (1.2) defines a self-adjoint operator which generates by Stone’s theorem the unitary time evolution $e^{-H_{N}t}$.
@@
$i\frac{\textnormal{d}}{\textnormal{d}t}\psi_{t}=H_{N}\psi_{t},\qquad\psi_{0}\in L^{2}(\Lambda,\textnormal{d}y)\otimes\mathcal{H}_{N}^{-}.$ (1.3)
Note that the ground state of the non-interacting Fermion system is well-known, non-degenerate and explicitly given by
@@
with a general state for the impurity particle i.e. the system is initially prepared in a state describing the ground state of the ideal Fermi gas which does not interact with the impurity particle.
@@
which is a consequence of Gauss’ counting algorithm.
1.2 Scaling regimes
@@
In the following, we present the choices of $\beta,\kappa$ and $\lambda$ we are aiming for, and discuss the physical meaning of the couplings.
@@
  • The coupling $\lambda$ models the coupling strength between the Fermi gas and the impurity. For small $\lambda$ we expect a decoupling between the gas and the impurity in the sense that the time evolution given by (1.3) does not entangle an initial state of product form $\phi\otimes\Omega_{0}$. Such a result was shown in [MP21] with $\beta,\kappa=1$ and $\lambda=k_{\textnormal{F}}^{-1/2}\sim N^{-\frac{1}{6}}$ in three dimensions and [JMPP17, JMP18] with $\beta=\kappa=\lambda=1$ in two dimensions.
  • The couplings $\beta$ and $\kappa$ determine the order of the kinetic energy of the impurity and Fermi gas respectively.
  • In addition, it is important to discuss on which time scales results hold. Define $\tau$ to be the relevant time scale for our statements, i.e. our statements $A(t)$ shall be considered for $t\in[0,T]$ with $T\in\mathcal{O}(\tau)$.
Our results hold for the following set of couplings
$S=\left\{(\beta,\kappa,\lambda,\tau):\beta\tau\in o(1),\ \kappa=\lambda,\ \lambda k_{\textnormal{F}}\tau\in\mathcal{O}(1)\right\}\subset\mathbb{R}_{\geq 0}^{4}.$ (1.8)
The physical interpretation can be better understood in the following specific cases:
Fermi time dynamics $\kappa=\lambda=1$, $\beta=k_{\textnormal{F}}^{-b}$ with $b>-1$, $\tau=k_{\textnormal{F}}^{-1}$
In this setting the kinetic energy of the Fermi gas and the interaction term are chosen independent of the the Fermi momentum $k_{\textnormal{F}}$ and, in particular of the particle number $N$. In return the considered time scale is $t\in\mathcal{O}(k_{\textnormal{F}}^{-1})$ and therefore short for large $k_{\textnormal{F}}$. Planck’s constant $\hbar$ is set to be $1$, the Fermion mass is of order 1. For convenience we set $m=1/2$ in this case. The mass $m_{\text{imp}}=\frac{1}{2}k_{\textnormal{F}}^{b}$ of the impurity particle might depend on $k_{\textnormal{F}}$ and represent for $b\in(-1,0)$ and large $k_{\textnormal{F}}$ a light, for $b>0$ a heavy or for $b=0$ an equally heavy impurity particle in comparison to the fermions. We remark that the results in [MP21] can be transferred to this setting of $\kappa=\lambda=1$, however, allowing only for shorter timescales of $\tau=k_{\textnormal{F}}^{-1-\delta}$ for a $\delta>0$. This seems compatible since we are not interested in a free time evolution of the impurity particle. The times $t\in\mathcal{O}(k_{\textnormal{F}}^{-1})$ are on the time scale of the fermions near the Fermi surface, which typically have a momentum of $k_{\textnormal{F}}$.
Heavy fermion regime $\kappa=\lambda=k_{\textnormal{F}}^{-1}$, $\beta=k_{\textnormal{F}}^{-b}$ with $b>0$, $\tau=1$
In this setting, we identify the couplings $\kappa$ and $\beta$ with the masses of the fermions and the impurity particle, respectively. More concretely, the fermion mass is $m=\frac{1}{2}k_{\textnormal{F}}$ and the mass of the impurity particle is $m_{\text{imp}}=\frac{1}{2}k_{\textnormal{F}}^{b}$ where the factor $\frac{1}{2}$ is chosen for convenience. Since we are interested in large $k_{\textnormal{F}}$, this case corresponds to a heavy Fermi gas setting where the impurity particle has a relative mass ratio of $\kappa/\beta$ depending on $b>0$. The coupling strength $\lambda$ between both impurity particle and gas is weak but much stronger than in the mean-field setting where one would introduce an averaging factor of $\lambda=N^{-1}\sim k_{\textnormal{F}}^{-3}$. Planck’s constant $\hbar$ is set to be $1$ in this case and is particularly independent of the number of gas particles.
Semiclassical regime $\kappa=\lambda=k_{\textnormal{F}}^{-2}$, $\beta=k_{\textnormal{F}}^{-b}$ with $b>1$, $\tau=1$
We can also identify the couplings with Planck’s constant $\hbar$ instead of setting it to $1$. In this case, we consider the time evolution governed by the Schrödinger equation
@@
$i\hbar\partial_{t}\psi_{t}=\left(\hbar^{2}\beta^{\prime}(-\Delta_{y})+\hbar^{2}\sum_{i=1}^{N}(-\Delta_{x_{i}})+\lambda\sum_{i=1}^{N}V(x_{i}-y)\right)\psi_{t}$ (1.9)
with $\hbar\coloneqq\kappa^{1/2}=\lambda^{1/2}=k_{\textnormal{F}}^{-1}$. The fermion mass is given by $m=1/2$ and mass of the impurity particle $m_{\text{imp}}=\beta^{\prime-1}=\frac{1}{2}k_{\textnormal{F}}^{b^{\prime}}$ with $b^{\prime}>-1$. The time evolution is therefore given by the operator $U_{t}=e^{-iH_{N}t/\hbar}$. Since the time scale $\tau=1$ is absorbed by the factor $t/\hbar$ is of order $k_{\textnormal{F}}$. [^1] Note that the effective Hamiltonian is expected to be of order $\hbar$ here such that the exponent of the time evolution operator is of order $1$. Since we are interested in $k_{\textnormal{F}}\gg 1$, the interpretation is that the system is considered in the semiclassical regime with $\hbar\ll 1$. Such a semi-classical regime with $\hbar\equiv k_{\textnormal{F}}^{-1}$ has been widely studied in the analysis of Fermi gases, as can be seen for example in [Ben22, Saf23].
From a technical point of view all settings are very similar and our results will apply to all described cases. We will focus our presentation on the setting with $\hbar\equiv 1$ and give some remarks how the results can be translated to the semiclassical setting.
@@ 2 Preliminaries
2.1 Second quantization
It is convenient to consider $L^{2}(\Lambda,\textnormal{d}y)\otimes\mathcal{H}_{N}^{-}$ as $N$-particle sector of $L^{2}(\Lambda)\otimes\mathcal{F}$ with the fermionic Fock space $\mathcal{F}$ constructed over $L^{2}(\Lambda)$. This way, we have access to the powerful formalism of second quantization with the fermionic creation operator $a_{p}^{\ast}$ creating a particle with momentum $p\in\mathbb{Z}^{3}$ and the annihilation operator $a_{p}$ annihilating a particle with momentum $p\in\mathbb{Z}^{3}$. Those operators satisfy the canonical anticommutation relations (CAR)
@@ 2.1 Second quantization
$\mathbb{H}=\underbrace{-\beta\Delta_{y}}_{\eqqcolon h_{0}}+\underbrace{\kappa\sum_{k\in\mathbb{Z}^{3}}|k|^{2}a_{k}^{\ast}a_{k}}_{\eqqcolon\mathbb{H}^{\textnormal{kin}}}+\underbrace{\lambda\sum_{k,p\in\mathbb{Z}^{3}}\hat{V}(k)e^{iky}a_{p}^{\ast}a_{p-k}}_{\eqqcolon\mathbb{V}}$ (2.2)
@@
2.2 Particle-hole transformation
In our analysis, a primary objective is to focus on excitations relative to the non-interacting Fermi ball. In particular, we want to use a description of our fermionic system in which the non-interacting Fermi ball $\Omega_{0}=\prod_{k\in B_{\textnormal{F}}}a^{\ast}_{k}\Omega$ is mapped to the vacuum. To achieve this, we employ the particle-hole transformation, which is a specific type of fermionic Bogoliubov transformation as creation operators are mapped to linear combinations of creation and annihilation operators while preserving the CAR. The particle-hole transformation is defined as the map $R:\mathcal{F}\to\mathcal{F}$ satisfying
$R^{\ast}a_{k}^{\ast}R\coloneqq\begin{cases}a_{k}^{\ast}&,k\in B_{\textnormal{F}}^{c}\\ a_{k}&,k\in B_{\textnormal{F}}\end{cases},\qquad R\Omega\coloneqq\Omega_{0}\ .$ (2.3)
@@ 2.2 Particle-hole transformation
$E_{N}^{\text{pw}}\coloneqq\kappa\sum_{k\in B_{\textnormal{F}}}|k|^{2}=\langle R\Omega,H_{N},R\Omega\rangle$ (2.5)
@@
$\mathbb{H}_{0}$ $\coloneqq R^{\ast}\mathbb{H}^{\textnormal{kin}}R-E_{N}^{\text{pw}}=\kappa\sum_{k\in\mathbb{Z}^{3}}|k|^{2}R^{\ast}a_{k}^{\ast}RR^{\ast}a_{k}R-E_{N}^{\text{pw}}$ (2.6)
$=\kappa\sum_{k\in B_{\textnormal{F}}}|k|^{2}a_{k}a_{k}^{\ast}+\kappa\sum_{k\in B_{\textnormal{F}}^{c}}|k|^{2}a_{k}^{\ast}a_{k}-E_{N}^{\text{pw}}$ (2.7)
$=\sum_{k\in\mathbb{Z}^{3}}e(k)a_{k}^{\ast}a_{k}\qquad\text{with }e(k)\coloneqq\begin{cases}\kappa|k|^{2}&,k\in B_{\textnormal{F}}^{c}\\ -\kappa|k|^{2}&,k\in B_{\textnormal{F}}.\end{cases}$ (2.8)
@@
$R^{\ast}\mathbb{H}R=-\beta\Delta_{y}+\mathbb{H}_{0}+b^{\ast}(\tilde{h})+b(\tilde{h})+\mathcal{E}$ (2.12)
with $\tilde{h}(k)\coloneqq\tilde{h}_{y}(k)\coloneqq\lambda\hat{V}(k)e^{iky}$ and $\mathcal{E}$ is given by the terms of (2.9b) since
@@
2.3 Almost-bosonic operators and patch decomposition
Our effective description of the microscopic system described by (1.2) will involve the emergence of almost-bosonic particles describing pair excitations of the Fermi ball. Those pair excitations will be delocalized over the Fermi surface in the sense that they correspond to a linear combination of two fermionic creation operators. As mentioned before the almost-bosonic pair operators which occur in this article coincide with the ones introduced in the series of seminal works [BNP+19, BNP+21a, BNP+21b, BPSS23] on the correlation energy of a weakly interacting Fermi gas. We give a brief introduction to the construction of those operators with the most relevant properties in the this subsection and in Section A.
@@ 2.3 Almost-bosonic operators and patch decomposition
$\Gamma^{\textnormal{nor}}\coloneqq\left\{(k_{1},k_{2},k_{3})\in\mathbb{Z}^{3}\cap\textnormal{supp}\hat{V}:k_{3}>0\vee k_{3}=0,k_{2}>0\vee k_{2}=k_{3}=0,k_{1}>0\right\}$ (2.14)
allowing the decomposition $\Gamma^{\textnormal{nor}}\sqcup(-\Gamma^{\textnormal{nor}})=\mathbb{Z}^{3}\cap\textnormal{supp}\hat{V}$.
@@
    Equivalently, we can also write the condition as $M=N^{c\delta}$ with $\delta\in(0,1/6)$ and $c\in(2,\frac{2}{3\delta}-2$). The lower bound on $M$ is needed to control the number of momenta inside each patch whereas the upper bound is needed to suppress Pauli’s principle. The choice of $\delta$ and $c$ will be taken later.
@@
    Define equal-area disjoint patches $p_{\alpha}$ centered around $\omega_{\alpha}\in S^{2}$ as follows
@@
    For given $k\in\Gamma^{\textnormal{nor}}$ define the index set $\mathcal{I}_{k}\coloneqq\mathcal{I}_{k}^{+}\cup\mathcal{I}_{k}^{-}$ consisting of north and south patch indices
@@
    $\mathcal{I}_{k}^{-}$$\coloneqq\{\alpha\in\{1,\ldots,M\}\mid k\cdot\hat{\omega}_{\alpha}\leq-N^{-\delta}\}.$ (2.17)
    A thin strip around the equator has to be excluded since the number of momenta per equator patch is too small. Note that $\delta>0$ coincides with the parameter in step 1.
@@
    $c_{\alpha}^{\ast}(k)\coloneqq b_{\alpha}^{\ast}(\pm k),\ \ n_{\alpha}(k)\coloneqq m_{\alpha}(\pm k)\iff\alpha\in\mathcal{I}_{k}^{\pm}$ (2.18)
@@
    $b_{\alpha}^{\ast}(k)\coloneqq\frac{1}{m_{\alpha}(k)}\sum_{p\in B_{F}^{c}\cap B_{\alpha},p-k\in B_{F}\cap B_{\alpha}}a_{p}^{\ast}a_{p-k}^{\ast}$ (2.19)

### v2

@@ Abstract
We study the quantum dynamics of a homogeneous ideal Fermi gas coupled to an impurity particle on a three-dimensional box with periodic boundary condition. For large Fermi momentum $k_{\textnormal{F}}$, we prove that the effective dynamics is generated by a Fröhlich-type polaron Hamiltonian, which linearly couples the impurity particle to an almost-bosonic excitation field. Moreover, we prove that the effective dynamics can be approximated by an explicit coupled coherent state. Our method is applicable to a range of interaction couplings, in particular including interaction couplings of order 1 and time scales of the order $k_{\textnormal{F}}^{-1}$.
@@ 1 Introduction
The study of impurities in quantum gases has garnered considerable attention due to its relevance in various physical contexts, ranging from solid-state physics to cold atom experiments. In this context quasi-particles such as polarons stand as intriguing entities emerging from the interaction of a single impurity particle with a surrounding medium. The concept of a polaron, originally introduced by Lev Landau to study the motion of an electron in a dielectric crystal [Lan65], most famously emerges from the celebrated Fröhlich Hamiltonian in second quantization formalism describing electron-phonon interactions [Frö54]. Subsequently, the polaron concept was extended to all kind of surrounding media including Bose and Fermi gases. The formation conditions and properties of polarons are believed to play a central role to understand the transport properties and the effective mass of impurities within the host material.
In this article, we study with mathematical rigor the dynamics of an impurity particle immersed in a dense gas of fermions as surrounding medium. Interactions between fermions are neglected and we assume that the impurity particle interacts with the Fermi gas via a finite-range potential in momentum space. The initial state $\psi$ of the system is a product state between the impurity state and a filled Fermi ball. This mathematical framework finds resonance with recent experimental and theoretical advancements in the study of ultracold atoms [SWSZ09, KSN+12, CJL+15]. We emphasize that our assumption on the interaction potential aligns with the growing interest among experimentalists and theorists in cases where the impurity is charged, and hence the interaction potential is long-ranged in position space [CCGB22, MJ24]. We show that the effective dynamics of the system is governed by a Fröhlich-type Hamiltonian, which linearly couples the impurity particle to an almost-bosonic excitation field. More specifically, the excitations relative to the filled Fermi ball are up to a constant described by the Hamiltonian
@@
Subsequently, we show that the effective time evolved state can be up to a phase factor approximated by a time-dependent coupled coherent state $W(\eta_{t})\phi\otimes\Omega$ where $W$ is the Weyl operator of the excitation field which is simply parameterized by a function $\eta_{t}$ and $\Omega$ is the Fock space vacuum. An explicit expression for $\eta_{t}$ is derived which allows for determining the number of collective excitations over time. We believe that such quantities are particularly helpful to gain deeper insights into the formation process of quasi-particles as studied in experiments such as [CJL+16]. Eventually, we show that the linear coupling term $\Phi(h_{y})$ cannot be omitted in the effective description but adds a leading order effect to the effective dynamics in the setting of interaction couplings of order 1.
Our results hold for a variety of time scales and couplings, describing different interaction strengths and mass ratios, which will be specified in the subsequent section. We note that the same microscopic model has been studied in [JMPP17, JMP18, MP21] but with very specific choices of couplings different from ours, leading to an effective decoupling of impurity and gas.
@@ 1.1 The microscopic model
We consider an impurity particle interacting with $N$ spinless fermions on a 3-dimensional box with periodic boundaries described by $\Lambda\coloneqq\mathbb{T}^{3}\coloneqq\mathbb{R}^{3}/(2\pi\mathbb{Z}^{3})$. The system is described by a state in the Hilbert space $L^{2}(\Lambda,\textnormal{d}y)\otimes\mathcal{H}_{N}^{-}$ with $\mathcal{H}_{N}^{-}=L^{2}(\Lambda)^{\wedge N}$ where $y$ is the coordinate of the impurity particle and $\{x_{i}\}_{i=1,\ldots,N}$ are the coordinates of the fermions. The Hamiltonian for our main model of the system is given by
$H_{N}=-\beta\Delta_{y}+\sum_{i=1}^{N}(-\Delta_{x_{i}})+\lambda\sum_{i=1}^{N}V(x_{i}-y)$ (1.2)
and parameterized by $\beta,\lambda>0$. Note that the different parts of the Hamiltonian on different tensor components of our Hilbert space writing, i.e. we used the short-hand notation writing, for example, $-\Delta_{y}\coloneqq-\Delta_{y}\otimes 1$ for the Laplacian acting on the impurity particle. The interaction $V$ is assumed to have a Fourier transform $\hat{V}$ with compact support satisfying $\hat{V}(k)=\hat{V}(-k)$ and $\hat{V}(k)\geq 0$ for all $k\in\mathbb{Z}^{3}$. It is well-known that under this assumption the Hamiltonian (1.2) defines a self-adjoint operator which generates by Stone’s theorem the unitary time evolution $e^{-H_{N}t}$.
@@
$i\hbar\frac{\textnormal{d}}{\textnormal{d}t}\psi_{t}=H_{N}\psi_{t},\qquad\psi_{0}\in L^{2}(\Lambda,\textnormal{d}y)\otimes\mathcal{H}_{N}^{-}.$ (1.3)
where $\hbar$ is the reduced Planck constant. For mathematical convenience we will always set $\hbar=1$ unless explicitly stated otherwise.
Note that the filled Fermi ball is a ground state for the non-interacting Fermion system. It is non-degenerate and explicitly given by
@@
with a general state $\phi$ for the impurity particle i.e., the system is initially prepared in a state describing a ground state of the ideal Fermi gas which does not interact with the impurity particle.
@@
which is a consequence of Gauss’ counting argument.
1.2 Relevant parameters and time scales
@@
In the following, we present the ranges of $\beta$ and $\lambda$ we are aiming for, and discuss the physical meaning of the parameters. In addition, it is important to discuss on which time scales our results hold.
  • The parameter $\beta$ determines the mass ratio of the impurity and Fermi gas particles such that $\beta\ll 1$ corresponds to a relatively heavy, $\beta\gg 1$ to a relatively light impurity particle. The case of equal masses corresponds to $\beta=1$. Our work requires the restriction $\beta\in o(k_{\textnormal{F}})$ see Remark 3.7. Since $k_{\textnormal{F}}$ tends to infinity, this allows to include all three cases.
@@
  • The parameter $\lambda\in[k_{\textnormal{F}}^{-1/6},1]$ models the coupling strength between the Fermi gas and the impurity. For small $\lambda$ we expect a decoupling between the gas and the impurity in the sense that the time evolution given by (1.3) does not entangle an initial state of product form $\phi\otimes\Omega_{0}$. Such a result was shown in [MP21] with $\beta=1$ and $\lambda=k_{\textnormal{F}}^{-1/2}$ in three dimensions and [JMPP17, JMP18] with $\beta=\lambda=1$ in two dimensions. Note that our interaction coupling $\lambda$ is much larger than in [MP21] and in particular, we are able to include $\lambda=1$. As will be discussed in Remark 3.10, it is necessary to include excitations of the Fermi gas into the effective description for $\lambda=1$.
  • Our approximations will apply for times $t\in\mathcal{O}(k_{\textnormal{F}}^{-1}\lambda^{-1})$, so that a weaker coupling strength $\lambda$ corresponds to slightly larger times. The times $t\in\mathcal{O}(k_{\textnormal{F}}^{-1})$ are on the time scale of the fermions near the Fermi surface, which have an approximate momentum of $k_{\textnormal{F}}$ and thus travel a distance of order 1 in this time. We are therefore able to enter a time scale where the impurity particle can resolve the motion of the fermions. We remark that the results in [MP21] can be transferred to this setting of $\lambda=1$, however, allowing only for shorter time scales of $t\in o(k_{\textnormal{F}}^{-1})$.
We remark that results obtained in the previously mentioned parameter range remain valid under re-scaling by multiplying the Hamiltonian by an overall factor and absorbing it by re-scaling the time variable. Although we will stick to the above choices, it might still be helpful for the reader to see some connections to other scaling regimes by re-scaling. We present two of them briefly:
  • Time scales of order 1 are obtained by multiplying the Hamiltonian by an overall factor of $k_{\textnormal{F}}^{-1}$. In this case, the factor $k_{\textnormal{F}}^{-1}$ appears as new parameters in front of the kinetic energy of the Fermi gas corresponding to a heavy fermion regime.
  • In recent years the so-called semiclassical regime has been widely studied in the analysis of dense Fermi gases, as can be seen for example in [Ben22, Saf23]. This regime is associated with identifying $\hbar\coloneqq k_{\textnormal{F}}^{-1}$ as small parameter instead of setting $\hbar=1$. It is achieved by multiplying the Hamiltonian by an overall factor of $k_{\textnormal{F}}^{-2}$ and absorbing $k_{\textnormal{F}}^{-1}$ in the time variable. For this new time scale, our theorem makes a statetment for times of order 1. The re-scaled Schrödinger equation takes the form of
@@
    $i\hbar\partial_{t}\psi_{t}=\left(\hbar^{2}\beta(-\Delta_{y})+\hbar^{2}\sum_{i=1}^{N}(-\Delta_{x_{i}})+\lambda k_{\textnormal{F}}^{-2}\sum_{i=1}^{N}V(x_{i}-y)\right)\psi_{t}$ (1.8)
    One identifies $\lambda k_{\textnormal{F}}^{-2}\in[k_{\textnormal{F}}^{-13/6},k_{\textnormal{F}}^{-2}]$ as re-scaled interaction coupling.
@@ 2 Preliminaries
Second quantization
It is convenient to consider $L^{2}(\Lambda,\textnormal{d}y)\otimes\mathcal{H}_{N}^{-}$ as the $N$-particle sector of $L^{2}(\Lambda)\otimes\mathcal{F}$ with the fermionic Fock space $\mathcal{F}$ constructed over $L^{2}(\Lambda)$. This way, we have access to the powerful formalism of second quantization with the fermionic creation operator $a_{p}^{\ast}$ creating a particle with momentum $p\in\mathbb{Z}^{3}$ and the annihilation operator $a_{p}$ annihilating a particle with momentum $p\in\mathbb{Z}^{3}$. Those operators satisfy the canonical anticommutation relations (CAR)
@@
$\mathbb{H}=\underbrace{-\beta\Delta_{y}}_{\eqqcolon h_{0}}+\underbrace{\sum_{k\in\mathbb{Z}^{3}}|k|^{2}a_{k}^{\ast}a_{k}}_{\eqqcolon\mathbb{H}^{\textnormal{kin}}}+\underbrace{\lambda\sum_{k,p\in\mathbb{Z}^{3}}\hat{V}(k)e^{iky}a_{p}^{\ast}a_{p-k}}_{\eqqcolon\mathbb{V}}$ (2.2)
@@
Particle-hole transformation
In our analysis, the primary objective is to focus on excitations relative to the non-interacting Fermi ball. In particular, we want to use a description of our fermionic system in which the non-interacting Fermi ball $\Omega_{0}=\prod_{k\in B_{\textnormal{F}}}a^{\ast}_{k}\Omega$ is mapped to the vacuum. To achieve this, we employ the particle-hole transformation, which is a specific type of fermionic Bogoliubov transformation as creation operators are mapped to linear combinations of creation and annihilation operators while preserving the CAR. The particle-hole transformation is defined as the map $R:\mathcal{F}\to\mathcal{F}$ satisfying
$R^{\ast}a_{k}^{\ast}R\coloneqq\begin{cases}a_{k}^{\ast}&\text{if }k\in B_{\textnormal{F}}^{c},\\ a_{k}&\text{if }k\in B_{\textnormal{F}}\end{cases}\quad\text{and}\quad R\Omega\coloneqq\Omega_{0}\ .$ (2.3)
@@
$E_{N}^{\text{pw}}\coloneqq\sum_{k\in B_{\textnormal{F}}}|k|^{2}=\langle R\Omega,H_{N},R\Omega\rangle$ (2.5)
@@
$\mathbb{H}_{0}$ $\coloneqq R^{\ast}\mathbb{H}^{\textnormal{kin}}R-E_{N}^{\text{pw}}=\sum_{k\in\mathbb{Z}^{3}}|k|^{2}R^{\ast}a_{k}^{\ast}RR^{\ast}a_{k}R-E_{N}^{\text{pw}}$ (2.6)
$=\sum_{k\in B_{\textnormal{F}}}|k|^{2}a_{k}a_{k}^{\ast}+\sum_{k\in B_{\textnormal{F}}^{c}}|k|^{2}a_{k}^{\ast}a_{k}-E_{N}^{\text{pw}}$ (2.7)
$=\sum_{k\in\mathbb{Z}^{3}}e(k)a_{k}^{\ast}a_{k}\qquad\text{with }e(k)\coloneqq\begin{cases}|k|^{2}&\text{if }k\in B_{\textnormal{F}}^{c},\\ -|k|^{2}&\text{if }k\in B_{\textnormal{F}}.\end{cases}$ (2.8)
@@
$R^{\ast}\mathbb{H}R=-\beta\Delta_{y}+\mathbb{H}_{0}+b^{\ast}(\tilde{h}_{y})+b(\tilde{h}_{y})+\mathcal{E}$ (2.12)
with $\tilde{h}_{y}(k)\coloneqq\lambda\hat{V}(k)e^{iky}$ and $\mathcal{E}$ is given by the terms of (2.9b) since
@@
Almost-bosonic operators and patch decomposition
Our effective description of the microscopic system described by (1.2) will involve the emergence of almost-bosonic particles describing pair excitations of the Fermi ball. Those pair excitations will be delocalized over the Fermi surface in the sense that they correspond to a linear combination of pairs of fermionic operators. As mentioned before the almost-bosonic pair operators which occur in this article coincide with the ones introduced in the series of seminal works [BNP+19, BNP+21a, BNP+21b, BPSS23] on the correlation energy of a weakly interacting Fermi gas. We give a brief introduction to the construction of those operators with the most relevant properties in this subsection and in Section A.
@@
$\Gamma\coloneqq\left\{(k_{1},k_{2},k_{3})\in\mathbb{Z}^{3}\cap\textnormal{supp}\hat{V}:(k_{3}>0\vee k_{3}=0),(k_{2}>0\vee k_{2}=k_{3}=0),k_{1}>0\right\}$ (2.14)
allowing the decomposition $\Gamma\cup(-\Gamma)=\mathbb{Z}^{3}\cap\textnormal{supp}\hat{V}$.
@@
    The lower bound on $M$ is needed to control the number of momenta inside each patch whereas the upper bound is needed to suppress Pauli’s principle. The choice of $\delta$ and $c$ will be taken later.
@@
    Define equal-area disjoint patches $p_{\alpha}$ as follows
@@
    Define $\omega_{\alpha}\in S^{2}$ as centers of the patch $p_{\alpha}$. A useful graphical sketch of this patch construction is given in Figure 1 of [BNP+19].
@@
    For given $k\in\Gamma$ define the set of north and south patch indices
@@
    $\mathcal{I}_{k}^{-}$$\coloneqq\{\alpha\in\{1,\ldots,M\}\mid k\cdot\hat{\omega}_{\alpha}\leq-N^{-\delta}\}$ (2.17)
    and $\mathcal{I}_{k}\coloneqq\mathcal{I}_{k}^{+}\cup\mathcal{I}_{k}^{-}$. This has the effect of excluding a strip around the equator of the Fermi ball, where the number of momenta per patch may become too small. Note that $\delta>0$ coincides with the parameter in step 1.
@@
    $c_{\alpha}^{\ast}(k)\coloneqq\begin{cases}b_{\alpha}^{\ast}(+k)&\text{if }\alpha\in\mathcal{I}_{k}^{+},\\ b_{\alpha}^{\ast}(-k)&\text{if }\alpha\in\mathcal{I}_{k}^{-}.\end{cases},\ \ n_{\alpha}(k)\coloneqq\begin{cases}m_{\alpha}(k)&\text{if }\alpha\in\mathcal{I}_{k}^{+},\\ m_{\alpha}(-k)&\text{if }\alpha\in\mathcal{I}_{k}^{-}.\end{cases}$ (2.18)
@@
    $b_{\alpha}^{\ast}(k)\coloneqq\frac{1}{m_{\alpha}(k)}\sum_{\begin{subarray}{c}p\in B_{F}^{c}\cap B_{\alpha},\\ p-k\in B_{F}\cap B_{\alpha}\end{subarray}}a_{p}^{\ast}a_{p-k}^{\ast},\quad m_{\alpha}(k)^{2}\coloneqq\sum_{\begin{subarray}{c}p\in B_{F}^{c}\cap B_{\alpha},\\ p-k\in B_{F}\cap B_{\alpha}\end{subarray}}1$ (2.19)